### v1

@@ 1.1. Background
Regarding the Dirichlet problem for the tangent Stokes system (S) (or any of its variants) on compact $C^{\infty}$ Riemannian manifolds $M$ with boundary, we mention Taylor’s book [72, Ch. 17, Ap. A] for the $\mathrm{L}^{2}$-based regularity theory for the Hodge Laplacian $\Delta_{H}$ and $M$ with smooth boundary. For the case of the deformation Laplacian $2\operatorname{Def}^{*}\operatorname{Def}$ (the analogue of the surface diffusion operator acting on $1$-forms) and $M$ with $C^{1}$ or Lipschitz boundary, existence of solutions in Sobolev–Besov spaces has been addressed by, among others, Dindos, Mitrea and Taylor [48, 20]. The arguments in [72, 48, 20] are built upon Schwartz kernel representation of differential operators and the representation of the velocity solution in the form of layer potentials.
More recently, Olshanskii, Reusken and Zhiliakov’s [61, Lemma 2.1] provide $\mathbf{H}^{2}\times\mathrm{H}^{1}$-regularity for the case of the surface diffusion operator $\Delta_{S}$ and 2D compact $C^{2,1}$-regular surfaces without boundary embedded in $\mathrm{R}^{3}$. Its proof hinges on the ($d=2$)-only characterization $\Delta_{H}=\operatorname{\boldsymbol{\operatorname{curl}}}_{\Gamma}\operatorname{curl}_{\Gamma}+\nabla_{\Gamma}\operatorname{div}_{\Gamma}$ of the Hodge Laplacian and its simple relation with the surface diffusion operator to decouple the velocity and the pressure, and regularity results in Hodge theory [54, §7.4]. There seem to be no readily available analogue for “rougher” manifolds of class $C^{m}$ ($m\in\mathrm{N}$) of arbitrary dimension $d\geq 2$ and Dindos, Mitrea and Taylor’s approach is not easily adaptable to this less stringent context, nor the argument in [61] can be extended as is to dimensions $d\geq 3$.
@@
For intrinsic dimensions $d\leq 4$, the weak formulation for the tangent Navier–Stokes reads: given $\mathbf{f}\in(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}$ and $g\in\mathrm{L}_{\#}^{2}(\Gamma)$, find $(\mathbf{u},p)\in\mathbf{H}^{1}_{t}(\Gamma)\times\mathrm{L}_{\#}^{2}(\Gamma)$ such that
@@
the restriction on $d$ is so as to ensure that the integral $\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{v}$ makes sense as a consequence of Sobolev embeddings. For $d\geq 5$, following Temam [75, Chapter 2] for the stationary Navier–Stokes system in flat domains, the testing space for the velocity must be reduced to $\mathbf{H}^{1}_{t}(\Gamma)\cap\mathbf{L}^{d}_{t}(\Gamma)$.
We are unable to refer to any specific results concerning the existence of solutions of (1.6) for data $\mathbf{f}$ and $g$ and manifold $\Gamma$ of minimum regularity, although we notice that a relatively straightforward adaptation of Temam’s strategy, that is, using a Faedo–Galerkin with eigenfunctions of an appropriate (tangent) Stokes operator (see Section 5 below) will suffice (see Section 6.2 below). On the other hand, we refer to the previously mentioned work by Dindos and Mitrea [20] for the existence of solutions in the Sobolev–Besov sense of the tangent Navier–Stokes system in $C^{\infty}$ manifolds $M$ of dimensions $d\leq 4$ and $C^{1}$ (or Lipschitz) boundary.
@@ 1.2. Statements of the main results
Let $\Gamma$ be of class $C^{\max\{2,m\}}$ for some nonnegative integer $m$, let $1<p<\infty$ and let $p^{*}:=\frac{p}{p-1}$ be its conjugate exponent.
Our first two main results concern the well-posedness and higher regularity for problems (LB) and (VL) problems. We find it important to mention that in this paper we only resort to parametrizations when proving $\mathrm{W}^{m,p},m\geq 2$ regularity for problem (LB); which is achieved by means of the powerful Calderón-Zygmund theory in flat domains. All other arguments are intrinsic to $\Gamma$, of “variational” nature and rely on non-obvious but elementary manipulations. We tackle the higher Sobolev regularity for problem (VL) in three separate cases: $p=2$, $p\in(1,2)$ and $p\in(2,\infty)$, which are related to each other via a so-called “variational duality”; see Figure 1. As part of this program, we also obtain the $\mathrm{W}^{1,p}$-well-posedness of problem (LB) by applying the Banach–Nečas–Babuška theorem (following ideas by Simader and Sohr [70]); this hinges on a non-trivial inf-sup condition in $\mathrm{W}^{1,p}_{\#}(\Gamma)\times\mathrm{W}^{1,p^{*}}_{\#}(\Gamma)$ that follows from a dual characterization of the $\mathbf{L}^{p}$-norm, the bijectivity of the Bochner–Laplace operator from $\mathbf{W}^{2,p^{*}}_{t}(\Gamma)$ onto $\mathbf{L}^{p^{*}}_{t}(\Gamma)$ and compactness arguments. The analogous result for problem (VL) follows a careful but elementary rewrite of the required inf-sup condition in terms of the one for (LB). All of these arguments are the subject of Section 3.
Theorem 1.1 (Laplace–Beltrami).
Let $\Gamma$ be of class $C^{3}$ and let $p\in(1,\infty)$. Then, for every $f\in(\mathrm{W}^{1,p^{*}}_{\#}(\Gamma))^{\prime}$, there exists a unique $u\in\mathrm{W}^{1,p}_{\#}(\Gamma)$ such that
Moreover, there exists $C=C(\Gamma,p)>0$ such that
Furthermore, if $\Gamma$ is of class $C^{\max\{3,m+2\}}$, for some nonnegative integer $m$ and $f\in\mathrm{W}^{m,p}_{\#}(\Gamma)$, then the solution $u\in\mathrm{W}^{1,p}_{\#}(\Gamma)$ of (1.7) is in fact in $\mathrm{W}^{m+2,p}_{\#}(\Gamma)$ and satisfies the PDE in strong form: $\Delta_{\Gamma}u=f$ a.e. in $\Gamma$. Moreover, there is $C=C(\Gamma,p)>0$ such that
$\|u\|_{m+2,p;\Gamma}\leq C\|f\|_{m,p;\Gamma}.$
We stress that 1.1 is our best attempt to condense in a sole statement our results concerning the Laplace–Beltrami operator at the expense of being sharp. In particular, the $C^{3}$-regularity of $\Gamma$ seems excessive when compared with the Lipschitz regularity for $p=2$. We urge the interested reader to refer to Section 3 for our precise results.
@@
Let $\Gamma$ be of class $C^{3}$ and let $p\in(1,\infty)$. Then, for every $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$, there exists a unique $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ such that
@@
Moreover, there exists $C=C(\Gamma,p)>0$ such that
@@
Moreover, if $\Gamma$ is of class $C^{m+3}$ for some nonnegative integer $m$ and $\mathbf{f}\in\mathbf{W}^{m,p}_{t}(\Gamma)$, then the solution $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ of (1.8) is in fact in $\mathbf{W}^{m+2,p}_{t}(\Gamma)$ and strongly satisfies the PDE: $\Delta_{B}\mathbf{u}=\mathbf{f}$ a.e. in $\Gamma$. Moreover, there is $C=C(\Gamma,p)>0$ s.t.
@@
We prove the $\mathrm{L}^{p}$-based ($1<p<\infty$) well-posedness and higher regularity of the tangent Stokes problem in Section 4 by generalizing a trick of Olshanskii, Reusken and Zhiliakov for the proof of [61, Lemma 2.1]. We begin by establishing the $\mathbf{H}^{1}\times\mathrm{L}^{2}$-well-posedness as a consequence of the classical Babuška–Brezzi theory in Hilbert spaces—this includes proving a Poincaré-type inequality for the covariant gradient $\nabla_{\Gamma}$. In turn, by means of non-trivial commutator identities on tangential calculus and exploting the lack of boundary of $\Gamma$ (cf. Appendix C), we obtain a collection of inf-sup conditions required to establish the $\mathbf{W}^{1,p}\times\mathrm{L}^{p}$-well-posedness as a direct application of the generalized Babuška–Brezzi theory in reflexive Banach spaces (see Appendix A). More precisely, this is achieved by decoupling the velocity and pressure unknowns into a Laplace–Beltrami problem for the pressure (with a right-hand side depending mildly on the velocity) and a Bochner–Laplace problem for the velocity, so as to be able to use the previously developed $\mathrm{L}^{p}$-based ($1<p<\infty$) well-posedness and Sobolev regularity for the Laplace–Beltrami and Bochner–Laplace operators (cf. 1.1 and 1.2), and $\mathbf{H}^{1}\times\mathrm{L}^{2}$-well-posedness of the tangent Stokes problem. Finally, the same decoupling procedure, combined with the aforementioned properties of the (LB) and (VL) problems and the $\mathbf{W}^{1,p}\times\mathrm{L}^{p}$-well-posedness of the tangent Stokes problem yields the higher $\mathrm{L}^{p}$-based Sobolev regularity results.
@@
Let $\Gamma$ be of class $C^{3}$ and let $p\in(1,\infty)$. Then, for every $(\mathbf{f},g)\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\times\mathrm{L}^{p^{*}}_{\#}(\Gamma)$, there exists a unique pair $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ such that
@@
Moreover, there exists $C=C(\Gamma,p)>0$ such that
$\|\mathbf{u}\|_{1,p;\Gamma}+\|\uppi\|_{0,p;\Gamma}\leq C\left(\|\mathbf{f}\|_{(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}}+\|g\|_{0,p^{*};\Gamma}\right).$
Furthermore, if $\Gamma$ is of class $C^{m+3}$ for some integer $m\geq 0$ and $(\mathbf{f},g)\in\mathbf{W}^{m,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$, then the solution $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ of (1.9) is in fact in $\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and satisfies the PDE in strong form: $\Delta_{B}\mathbf{u}+\nabla_{\Gamma}\uppi=\mathbf{f},\quad\operatorname{div}_{\Gamma}\mathbf{u}=g$ a.e. in $\Gamma$. Moreover, there is $C=C(\Gamma,p)>0$ such that
@@
Following ideas from [20, §§6] and [75, Chapter 2, §1], we are able to show existence of solutions to the tangent incompressible Navier–Stokes equations for dimensions $d\in\{2,3,4\}$ and integrability parameter $p\geq 2$.
To that end, we first focus our attention to a “linearized” version, the so-called tangent Oseen equations, of the incompressible tangent Navier–Stokes equations. By means of the Fredholm alternative in reflexive Banach spaces [12, Theorem 6.6], we establish the well-posedness of the Oseen equations. This is the subject of Section 6.1.
In turn, the existence of solutions to the incompressible Navier–Stokes equations for $p=2$ follows from a standard Galerkin method in terms of eigenfunctions of the Stokes operators (cf. Section 5). For $p>2$, we increase the integrability of solutions as a consequence of the well-posedness of the Oseen equations. This is the content of Section 6.2.
Theorem 1.4 (existence of solutions of the incompressible tangent Navier–Stokes equations for $d=2,3,4$).
Let $d\in\{2,3,4\}$, $2\leq p<\infty$ and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. Assume $\Gamma$ is of class $C^{3}$. Then, for each $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ there exists a solution $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ of
@@
For arbitrary dimensions $d\geq 2$, the well-posedness of the tangent Navier–Stokes is obtained in Section 6.3 by using the Banach fixed-point theorem under a restricted range of $p$ and a smallness assumption on data.
@@
Moreover, there exists $C=C(\Gamma,p,\mathbf{f},g)$ s.t.
@@
We bootstrap the aforementioned well-posedness and higher-regularity theory for the tangent Stokes problem (cf. 1.3) to deal with the nonlinearity $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}$ of (NS), resulting in the following higher regularity result for the tangent Navier–Stokes equations.
@@
Assume $\Gamma$ is of class $C^{m+3}$ for some integer $m\geq 0$ and $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ is a solution of (1.9) with right-hand side $(\mathbf{f},g)\in\mathbf{W}^{m,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$. Then, $(\mathbf{u},\uppi)$ actually lies in $\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and satisfies the PDE in strong form: $\Delta_{B}\mathbf{u}+(\nabla_{\Gamma}\mathbf{u})\mathbf{u}+\nabla_{\Gamma}\uppi=\mathbf{f},\quad\operatorname{div}_{\Gamma}\mathbf{u}=g$ a.e. in $\Gamma$. Moreover, the following estimate holds
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq F_{m,p}\big{(}\|\mathbf{f}\|_{m,p,\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{1,p;\Gamma}\big{)},$
@@
As it has already been mentioned, other operators can be chosen in lieu of the Bochner–Laplace $\Delta_{B}$ operator, such as Surface Diffusion operator $\mathbf{P}\operatorname{div}_{\Gamma}2D_{\Gamma}$ and Hodge Laplacian $\Delta_{H}$, as the leading second-order differential operator in problems (VL), (S), (NS); their difference from the Bochner–Laplace operator is explicitly stated in (1.4). Our preference for $\Delta_{B}$ is mostly for theoretical convenience, as the resulting weak formulations do not need to be analyzed “away from” kernel spaces, e.g. [40, Section 4]. Using identities (1.4), we show in Section 7 that solutions (provided they exist) to the aforementioned problems with $\Delta_{S}$ or $\Delta_{H}$ in place of $\Delta_{B}$ also enjoy $\mathrm{L}^{p}$-based higher Sobolev regularity.
Extensions to compact manifolds with boundary
Even though all of our results are stated for compact manifolds of class $C^{m}$ without boundary, we expect for some of them to extend to the case of a $d$-dimensional manifold $\Gamma$ of class $C^{m}$ with a sufficiently regular boundary $\partial\Gamma$. In particular,
  • higher regularity for the Laplace–Beltrami (LB) problem: when expressing the problem in parametric domain, boundary regularity estimates would need to be used in addition to the interior regularity estimates present in this paper.
Since our proofs for the vector-valued problems (vector Laplace (VL), tangent Stokes (S), tangent Navier–Stokes (NS)) rely crucially on geometric identities that are only valid on closed manifolds (because of integration-by-parts formulas), these results would not directly transfer to the case of $\Gamma$ with a non-empty boundary. We would instead need to formulate the problems in parametric domain in order to:
  • obtain higher-regularity estimates for vector-Laplace problem (VL) as a consequence of the $\mathrm{L}^{p}$-theory for elliptic systems with non-constant coefficients in flat domains (see, e.g., [33, Chapter 7]).
  • derive an $\mathrm{L}^{2}$-based Sobolev regularity theory for the tangent Stokes problem (S), and employ ideas by Beirão da Veiga [6] and Berselli [9] combined with an ADN-type result [2] to handle the $\mathrm{L}^{p}$-based Sobolev regularity theory.
Well-posedness in $\mathbf{W}^{1,p}$ for the Bochner–Laplace operator and $\mathbf{W}^{1,p}\times\mathrm{L}^{p}$ for the tangent (Navier–)Stokes equations would most likely have to be carried out via a Simader–Sohr-type argument [70].
@@
Since some of the results that we derived appear in various forms in the literature, we would like to emphasize which results, to the best of our knowledge, can only be found in this paper. For $d$-dimensional compact manifolds without boundary of class $C^{m}$ ($m\in\mathrm{N}$, $m\geq 2$), we make the following contributions:
  • $\mathrm{L}^{p}$-based ($1<p<\infty)$ well-posedness of weak formulations for problems (1.1). Notice that this is not an immediate consequence from the $\mathrm{L}^{2}$-based well-posedness or $\mathrm{L}^{p}$-based higher regularity.
  • $\mathrm{L}^{2}$-based regularity for the tangent Stokes problem for dimensions $d\geq 2$. The only similar result that we know of is for $C^{2,1}$-surfaces of dimension $d=2$ (for the Hodge Laplacian $\Delta_{H}$ and surface diffusion operator $\Delta_{S}$) [61, Lemma 2.1].
  • $\mathrm{L}^{p}$-based ($1<p<\infty$) regularity for the tangent Stokes and Navier–Stokes problems.

### v2

@@ 1.1. Background
Regarding the Dirichlet problem for the tangent Stokes system (S) (or any of its variants) on compact $C^{\infty}$ Riemannian manifolds $M$ with smooth boundary, we mention Taylor’s book [60, Ch. 17, Ap. A] for the $\mathrm{L}^{2}$-based regularity theory for the Hodge Laplacian $\Delta_{H}$. For the case of the deformation Laplacian $2\operatorname{Def}^{*}\operatorname{Def}$ (the analogue of the surface diffusion operator acting on $1$-forms) and smooth $M$ with $C^{1}$ or Lipschitz boundary, existence of solutions in Sobolev–Besov spaces has been addressed by, among others, Dindos, Mitrea and Taylor [42, 19]. The arguments in [60, 42, 19] are built upon Schwartz kernel representation of differential operators and the representation of the velocity solution in the form of layer potentials.
More recently, Olshanskii, Reusken and Zhiliakov’s [53, Lemma 2.1] provide $\mathbf{H}^{2}\times\mathrm{H}^{1}$-regularity for the case of the surface diffusion operator $\Delta_{S}$ and 2D compact $C^{2,1}$-regular surfaces without boundary embedded in $\mathrm{R}^{3}$. Its proof hinges on the ($d=2$)-only characterization $\Delta_{H}=\operatorname{\boldsymbol{\operatorname{curl}}}_{\Gamma}\operatorname{curl}_{\Gamma}+\nabla_{\Gamma}\operatorname{div}_{\Gamma}$ of the Hodge Laplacian and its simple relation with the surface diffusion operator to decouple velocity and pressure, and regularity results in Hodge theory [48, §7.4]. There seem to be no readily available analogue for “rougher” manifolds of class $C^{m}$ or $C^{m-1,1}$ ($m\in\mathrm{N}$) of arbitrary dimension $d\geq 2$ and Dindos, Mitrea and Taylor’s approach is not easily adaptable to this less stringent regularity context, nor can the argument in [53] be extended as is to dimensions $d\geq 3$.
@@
For intrinsic dimensions $d\leq 4$, the $\mathrm{L}^{2}$-based weak formulation for the tangent Navier–Stokes reads: given $\mathbf{f}\in(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}$ and $g\in\mathrm{L}_{\#}^{2}(\Gamma)$, find $(\mathbf{u},p)\in\mathbf{H}^{1}_{t}(\Gamma)\times\mathrm{L}_{\#}^{2}(\Gamma)$ such that
@@
the restriction $d\leq 4$ ensures that the integral $\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\mathbf{u}\cdot\mathbf{v}$ makes sense as a consequence of Sobolev embeddings.
We are unable to refer to any specific results concerning the existence of solutions of (1.5) for data $\mathbf{f}$ and $g$ and manifold $\Gamma$ of minimal regularity, although we notice that a relatively straightforward adaptation of Temam’s strategy, that is, using a Faedo–Galerkin method with eigenfunctions of an appropriate (tangent) Stokes operator (see Section 4.4 below) will suffice, as attested in Section 5.2 below. On the other hand, we refer to the aforementioned work by Dindos and Mitrea [19] for the existence of solutions in the Sobolev–Besov sense of the tangent Navier–Stokes system on $C^{\infty}$ manifolds of dimensions $d\leq 4$ and $C^{1}$ (or Lipschitz) boundary.
@@ 1.2. Statements of the main results
We now collect the main results of this paper, and briefly comment on their nuances and connections. The following result for the Bochner-Laplace operator (VL) is derived in Section 3 and hinges on a Poincaré inequality for the covariant gradient $\nabla_{\Gamma}$ proved in 3.3.
@@
Let $\Gamma$ be of class $C^{2}$ and let $p\in(1,\infty)$. Then, for every $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$, there exists a unique $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ such that
@@
Moreover, there exists a positive constant $C$, depending only on $\Gamma$ and $p$, such that
@@
Furthermore, if $\Gamma$ is of class $C^{m+2,1}$ for some nonnegative integer $m$ and $\mathbf{f}\in\mathbf{W}^{m,p}_{t}(\Gamma)$, then the solution $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ of (1.6) is in fact in $\mathbf{W}^{m+2,p}_{t}(\Gamma)$ and strongly satisfies the PDE: $\Delta_{B}\mathbf{u}=\mathbf{f}$ a.e. on $\Gamma$. In addition, there is a constant $C=C(\Gamma,p)>0$ such that
@@
We point out that the required regularity of $\Gamma$ in 1.1 is one unit higher than in [5, Theorems 1.1 and 1.2] for the scalar Laplace-Beltrami operator. This is due to the fact that the vector-valued solution $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ is tangent a.e. to $\Gamma$, whence $\mathbf{u}$ is a linear combination of tangent vectors to $\Gamma$ that already consume one derivative of the parametrization of $\Gamma$. This observation applies to all statements below that rely on 1.1.
We prove the $\mathrm{L}^{p}$-based ($1<p<\infty$) well-posedness and higher regularity of the tangent Stokes problem (S) in Section 4 by generalizing a trick of Olshanskii, Reusken and Zhiliakov for the proof of [53, Lemma 2.1].
@@
Let $\Gamma$ be of class $C^{2}$ and let $p\in(1,\infty)$. Then, for every $(\mathbf{f},g)\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\times\mathrm{L}^{p}_{\#}(\Gamma)$, there exists a unique pair $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ such that
@@
Moreover, there exists a positive constant $C$, depending only on $\Gamma$ and $p$, such that
Furthermore, if $\Gamma$ is of class $C^{m+2,1}$ for some nonnegative integer $m$ and $(\mathbf{f},g)\in\mathbf{W}^{m,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$, then the solution $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ is in fact in $\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and satisfies the PDEs in strong form: $\Delta_{B}\mathbf{u}+\nabla_{\Gamma}\uppi=\mathbf{f},\quad\operatorname{div}_{\Gamma}\mathbf{u}=g$ a.e. on $\Gamma$. In addition, there is a constant $C=C(\Gamma,p)>0$ such that
@@
The main argument goes as follows. We begin by establishing the $\mathbf{H}^{1}\times\mathrm{L}^{2}$-well-posedness as a consequence of the classical Babuška–Brezzi theory in Hilbert spaces—which hinges on 3.3 (Poincaré inequality). In turn, by means of non-trivial commutator identities on tangential calculus and exploiting the lack of boundary of $\Gamma$ (cf. Appendix B), we obtain a collection of inf-sup conditions required to establish the $\mathbf{W}^{1,p}\times\mathrm{L}^{p}$-well-posedness as a direct application of the generalized Babuška–Brezzi theory in reflexive Banach spaces (see Appendix A). More precisely, this is achieved by decoupling the velocity and pressure unknowns into a Laplace–Beltrami problem for the pressure (with a right-hand side depending mildly on the velocity) and a Bochner–Laplace problem for the velocity, so as to be able to use the previously developed $\mathrm{L}^{p}$-based ($1<p<\infty$) well-posedness and Sobolev regularity for the Laplace–Beltrami (cf. [5]) and Bochner–Laplace operators (cf. 1.1), and the $\mathbf{H}^{1}\times\mathrm{L}^{2}$-well-posedness of the tangent Stokes problem. Finally, the same decoupling procedure, combined with the aforementioned properties of the (LB) and (VL) problems and the $\mathbf{W}^{1,p}\times\mathrm{L}^{p}$-well-posedness of the tangent Stokes problem yields the higher $\mathrm{L}^{p}$-based Sobolev regularity results.
Following ideas from [19, §§6] and [63, Chapter 2, §1], we are able to show existence of solutions to the tangent incompressible Navier–Stokes equations (NS) for dimensions $d\in\{2,3,4\}$ and integrability parameter $p\in[2,\infty)$. To that end, we first focus our attention to a “linearized” version of (NS), the so-called tangent Oseen equations. By means of the Fredholm alternative in reflexive Banach spaces [12, Theorem 6.6], we establish the well-posedness of the Oseen equations. This is the subject of Section 5.1.
On the other hand, the existence of solutions to (NS) for $p=2$ follows from a standard Galerkin method in terms of eigenfunctions of the Stokes operators (cf. Section 4.4). For $p\in(2,\infty)$, we increase the integrability of solutions as a consequence of the well-posedness of the Oseen equations. This is the content of Section 5.2.
Let $d\in\{2,3,4\}$ and $p\in[2,\infty)$. Assume $\Gamma$ is of class $C^{2}$. Then, for each $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ there exists a solution $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ of
@@
For arbitrary dimensions $d\geq 2$, the well-posedness of (NS) is obtained in Section 5.3 by using the Banach fixed-point theorem under a restricted range of $p$ and a smallness assumption on data.
@@
Moreover, there exists a constant $C=C(\Gamma,p,\mathbf{f},g)>0$ such that
@@
We bootstrap the aforementioned well-posedness and higher-regularity theory for the tangent Stokes problem (cf. 1.2) to deal with the nonlinearity $(\nabla_{\Gamma}\mathbf{u})\mathbf{u}$ of (NS), resulting in the following higher regularity result for the tangent Navier–Stokes equations (cf. 5.9).
@@
Assume $\Gamma$ is of class $C^{m+2,1}$ for some nonnegative integer $m\geq 0$. Let $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ be a solution of (1.7) with right-hand side $(\mathbf{f},g)\in\mathbf{W}^{m,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$. Then, $(\mathbf{u},\uppi)$ actually lies in $\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and satisfies the PDE in strong form: $\Delta_{B}\mathbf{u}+(\nabla_{\Gamma}\mathbf{u})\mathbf{u}+\nabla_{\Gamma}\uppi=\mathbf{f}$, $\operatorname{div}_{\Gamma}\mathbf{u}=g$ a.e. on $\Gamma$. Moreover, the following estimate holds
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq F_{m,p}\big(\|\mathbf{f}\|_{m,p;\Gamma},\|g\|_{m+1,p;\Gamma},\|\mathbf{u}\|_{1,p;\Gamma}\big),$
where $F_{m,p}:\mathrm{R}^{3}\rightarrow[0,\infty)$ is a smooth nonlinear function, strictly increasing in each of its three components, and satisfies $F_{m,p}(0,0,0)=0$.
@@
As it has already been mentioned, other operators can be chosen in lieu of the Bochner–Laplace $\Delta_{B}$ operator, such as the Surface Diffusion operator $\Delta_{S}=\mathbf{P}\operatorname{div}_{\Gamma}2D_{\Gamma}$ and the Hodge Laplacian $\Delta_{H}$, as the leading second-order differential operator in problems (VL), (S) and (NS); their difference from the Bochner–Laplace operator is explicitly stated in (1.3). Our preference for $\Delta_{B}$ is mostly for theoretical convenience, as the resulting weak formulations do not need to be analyzed “away from” kernel spaces, e.g. [36, Section 4]. Using identities (1.3), we show in Section 6 that solutions (provided they exist) to the aforementioned problems with $\Delta_{S}$ or $\Delta_{H}$ in place of $\Delta_{B}$ also enjoy $\mathrm{L}^{p}$-based higher Sobolev regularity similar to $\Delta_{B}$.
@@
Since some of the results that we derived appear in various forms in the literature, we would like to emphasize which results, to the best of our knowledge, can only be found in this paper. For $d$-dimensional compact manifolds $\Gamma$ without boundary, we make the following contributions:
  1. $\bullet$
    $\mathrm{L}^{p}$-based well-posedness of weak formulations for problems (1.1) with minimal regularity of data, namely, $\Gamma$ of class $C^{2}$, $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ and $g\in\mathrm{L}^{p}_{\#}(\Gamma)$. Notice that this is not an immediate consequence of the $\mathrm{L}^{2}$-based well-posedness or $\mathrm{L}^{p}$-based higher regularity.
  2. $\bullet$
    $\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ regularity ($m\geq 0$, $1<p<\infty$) for the tangent Stokes problem for dimensions $d\geq 2$ provided $\Gamma$ is of class $C^{m+2,1}$. The only similar result that we know of is for $C^{2,1}$-surfaces of dimension $d=2$, $m=0$ and $p=2$ (for the Hodge Laplacian $\Delta_{H}$ and surface diffusion operator $\Delta_{S}$) [53, Lemma 2.1].
  3. $\bullet$
    $\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ regularity $(m\geq 0)$ for the Navier–Stokes problem for $p\in(p_{0},\infty)$ with $p_{0}=p_{0}(d)$ sufficiently far from $1$, provided $\Gamma$ is of class $C^{m+2,1}$.